UT Nov 15 2021 - D/E Night



Night Info

  • Sunset 17:25
  • Twilight End 18:44
  • Moonset 02:30
  • Twilght Begins 05:24
  • Sunrise 06:44

Night Plan

  • Configure for LBTI at sunset
  • DONE LBTI Photometric Standards 18:30hr-20:30hr
    • Conditions: clear skies, seeing <1.2"-1.5" (must be able to close the loop and stabilize)
  • Reconfigure for MODS 20:30-20:45 hr
  • DONE MODS Science: M31 TOO - 20:45hr-21:45 hr
  • Reconfigure for LBCs 21:45hr-22:15hr
  • DONE LBCs/TMS testing - 3hrs - 22:15hr-01:15hr
    • any conditions (good seeing, no clouds optimal)
    • TMS Setup
      • ssh telescope@robs
      • cd /lbt/lbto/supportscripts/MetrologySupport
      • Check if reference file is copied over (J. Hill or O. Kuhn will verify)
      • Go to co-point field, run dofpia
      • AFTER dofpia is completed then run TMS
        • First Set reference: ./lbc_set_tms_offset.py -s 2
          • this will take a while before returning prompt
        • In MetrologySupport directory execute ./tms_loop.pl once we are configured and authorized
      • co-point with lbcrangebal, then go to field of choice
      • Start running scripts
      • CTRL-C before running DOFPIA again - if run DOFPIA and TMS simultaneously, the two will "fight" - TMS will backout anything DOFPIA does!
  • Reconfigure from LBCs to bino-LUCIs 01:15hr-01:40hr
  • DONE Science: Arp 187 LUCI-Bino 01:40hr-03:40hr
  • Reconfigure to MODS-Bino 03:40hr-03:55hr
  • MODS Science: FF0937 - 04hr-twilight
  • Poor Weather Backup:
    • Arp 187 - proceed until Arp 187 sets (04:15)
    • Local Mergers until twilight


Staff:

  • OSA: D.G.Huerta
  • ISp: Mtn Staff
  • ISA: B. Rothberg, R.T.Gatto
  • Mtn Manager: M. Wagner

Night Log:

Closed Dome Checks:

  • LUCI software will be restarted and beta version of it run and beta version of RTD run on OBS5 tonight
  • UT 23:22 (-1 day) - waking MODS
  • UT 23:28 (-1 day) - simSnaps
    • OK - dust doughnut in MODS1 both red and blue
  • UT 23:37 (-1 day) - powering up LBCs
  • UT 23:41 (-1 day) - taking test biases
    • OK
  • UT 00:12 - no go on the LUCI testing - RTD changes on OBS5 - G.Eyechaner needs to arrange with IT about updating/changing software on the machine - postponed for another E/D night
  • UT 00:13 - bringing up the LUCI GUIS - init all and test darks
    • OK

Opening Dome:

LBTI
  • UT 00:15 - opening enclosure for LBTI
  • UT 03:40 - handover from LBTI to LBTO for MODS reconfigure
MODS-Bino Science - RMWagner
  • Start with Feige 110
  • UT 03:43 - puttng MODS into observng mode, homing AGWs
  • UT 03:52 - reconfig complete
  • UT 03:55 - acquire preset for pointing check
  • UT 03:59 - collimation preset
  • UT 04:01 - sending preset to Feige 110 (SpecStd)
    • preset hung on MODS-1 - CTRLC'd on M1
    • MODS2 hung on CLEARSTARS
    • CTRL C'd on M2 -- still went to target anyways
  • UT 04:04 - resending preset after Guiding started and preset completed
    • worked
  • Feige 110
    • M2 offsetxy 4.155 9.466 rel
    • M1 offsetxy -0.869 12.434 rel
    • tweaking M2 with modsAlign -r offsetxy -0.399 -0.349 rel
    • tweaking M1 with modsAlign -r offsetxy 0.312 -0.134 rel
    • UT 04:11 - starting Spectra
    • UT 04:14 - MODS-1 timed out on the obs script - resending - config just hung
    • M2 will finish before M1
  • M31 Nova
    • UT 04:27 - sending preset, sent PA=-90 first, CTRLC'd and will resend PA=0
    • UT 04:31 - resnet with PA=0
    • thru-slit is 60sec, could change to 15sec
    • M1 offsetxy -0.242 10.619 rel
    • M2 offsetxy 4.125 8.489 rel
    • M1 offsetting -0.12"
    • M2 offsetting +0.12"
    • 30sec check thru-slit
    • UT 04:44 - starting spectra
    • UT 04:46 guiders show 1", WFS 0.8"-0.9"
    • UT 04:48 - seeing bubble up to 1.4" but it came back down to 1.1"-1.2"
    • UT 04:50 - seeing variable again
    • 120sec Full Frame - 3k: instconfig red imaging, select filter (sdss r), manually click IMCS,
    • UT 05:27 - Reconfigure for LBCs
LBC TMS Testing
  • 3 hours of testing
TMS Plan
  • We plan to:
    • Start with a previous reference (04 Nov, which does not use channels 10 & 11)
    • Field 1: Slew to a high elevation field and run active tms for a few cycles
    • Slew to the copointing field for this field (F1)
      • Run dofpia
      • Run the copointing script
      • Set a new reference (while the copointing script is running)
      • Copoint
    • Observe Field 1 for a while
    • Pause tmsloop
    • Take a set of extrafocal pupils at Field1
      • run addsph.py to add z4/z11
      • Play the OB, RB_fastextra4.ob
      • run remsph.py to remove z4/z11
    • Restart tmsloop and
    • Slew to a low elevation field and wait for a few TMS cycles
    • Puse tmsloop and take a set of extrafocal pupils at Field2 (see details above0
    • Observe this field for about 1 hr
    • Pause tmsloop and take a set of extrafocal pupils at Field2
    • Resume tmsloop and slew to high elevation field and wait a few TMS cycles.
TMS Summary
TMS was run in active mode for 2 hours with the reference set at 06:46. Two fields were observed, 2004 QB at low elevation (43-48 deg), and 2013 MW6 at high elevation, (85-74 deg). The conditions were pretty stable though I noticed that the ambient temperature dropped about a degree (7 -> 6 deg C) during the observations. TMS held collimation during these observations. There were a few things to note:
  • Pointing/copointing off with an old reference: To start, we used the reference from the last time TMS was used, Nov 04 UT. It did not use channels 10 or 11. This did help with the collimation --- for the first red channel pupil image of the night, only Z4 was outside the convergence criterion (estimated seeing 0.56") and for the first blue channel pupil image, Z6, Z11 and Z22 were outside of the convergence criteria (estimated seeing 0.81"). However, the two sides were obviously not copointed and the IE,CA and mirror corrections that lbcrangebal sent were large: dIE = 50". The preset on the right side was lost. I removed the instrument offsets and we restored the Nov 04 UT reference in order to try the same at low elevation. Again, the pointing and copointing was off and lbcrangebal sent dIE = -47". This time the preset was not lost and the result was a pointed/copointed telescope. (The opposite sign makes me wonder whether I really reset everything by removing instrument offsets & setting the Nov 04 UT reference anew).
  • There were 3 images with trails or jumps:
    • lbcr.20211115.080923 has a 30" trail from where the stars were on the previous image to I suppose where they should be on this image. The associated blue image is not trailed.
    • lbcb.20211115.084008 and lbcr.20211115.083810 shows about 3 jumps covering about 2.5".
  • Operationally, it is difficult to notice when TMS has settled, i.e. sent about 3 sets of good corrections. So I think some efficiency is lost in watching and overdoing it because of losing count.
  • Also operationally, robs froze up during execution of the last OB. We let the OB continue, but when we wanted to run dofpia and pause the active loop, we had to go in through an ssh to kill the process. robs finally came back to life when reconfiguring for LUCI. The freeze-up lasted almost 30 min.

TMS Log
05:44 Done reconfiguring

05:45 Starting tmsloop.py (having already copied ref files from 20211104). There were some errors: pose calculation failed continuously (15 times):

**************************
**=> [2021-11-15 05:48:26.324445UT]
**=> requesting a TMS vector over TCP...
**=> parsing tcp reply...
****=> parsing failed (probably failed measurement)
**=> requesting weather telemetry...
****=> tsx=7.22C, tdx=6.94C, p=697.50hPa, h=20.00%
***************************
* pose calculation failed *
***************************
****************************************************************
* pose calculation failed continuously (15) times, aborting... *
****************************************************************
Traceback (most recent call last):
  File "/lbt/lbto/supportscripts/MetrologySupport/lbc_tms.py", line 725, in <module>
    sx_x=float(result[0].strip("[(,)]"))
ValueError: could not convert string to float: '*'
BOTH.... Sleeping for 5 seconds - ^C to exit

05:45 Heejoo is restarting the TMS software.

05:55 Let TMS work a bit at zenith and then slewed to copointing field for first field.

06:01 Having waited at least 3 cycles, now pausing the TMS loop and starting dofpia
  • dofpia "timed out" - first images took a while to appear in /newdata. OK, though.
  • red: all but z4 within convergence criteria on 1st pupil images, finally in 4 iter red converged (glass is cooler than ambient but only by 0.1 C)
  • blue: z6, z11 and z22 are all outside of convergence criteria on 1st pupil image, but converged on 2nd pupil image (est_seeing 1" and z6 was -281 nm)
06:10 Sending copointing script

06:11 Setting a new reference: ./lbc_set_tms_offset.py -s 2 and starting the loop. (./tms_loop.pl)

06:12 Copointing. Wow! This is quite far off - and we lost the preset on the right side

COPOINTING: B=61054 R=61049
Pointing updates: delta_IE = 50.59", delta_CA = 8.04"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:    0.17    -1.07   -22.90    -3.69
              DX:   -0.93    -2.40   -51.54    19.86

06:24 I am removing the instrument offsets (in order to remove the copointing, plus it removed filter focus offset).

06:25 Slewing to a low elevation field while running TMS. The reference that had been set at the end of Nov 04 - did it throw things out of whack?

06:29 Putting back the Nov 04 reference. Now, at the low elevation (elev 38 deg) field, running tms_loop.pl for a few cycles.
  • DX: not enough channels present, get a few nans. Took about 4 iterations before we got valid vectors, not nans. Sent 3 TMS correction and stopped.
06:37 Running dofpia now. Converged in 2 iterations. Then I took the pair of exposures on the copointing star and copointed.
COPOINTING: B=64243 R=64239
Pointing updates: delta_IE = -46.94", delta_CA = 1.79"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:    0.51     0.28     6.05   -10.94
              DX:   -0.50    -1.95   -41.81    10.63

06:46 Large offsets, but we are copointed now (did not lose the preset). taking another pair of exposures to verify: 064658, 06455. These look much better. Setting a new refernece here ./lbc_set_tms_offset.py -s 2

06:50 Copointed again - very tiny offsets. Set a new reference.

COPOINTING: B=64658 R=64655
Pointing updates: delta_IE = -1.03", delta_CA = -0.36"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:    0.03     0.01     0.12    -0.59
              DX:   -0.03    -0.03    -0.59     0.59

06:51 Starting TMS loop (./tms_loop.pl)

06:52 Observing the field, 2004 QB (elev 43) at g and r. Coords are for 6, not 7 UT, but object moved only a few arcseconds 25" in RA and 4" in DEC in the hour. Back-face corrections are being applied, but they are very small as the temperature has been stable:

SX Back-face corrections:     -86.50 Z4 (nm) or  -0.003278 dZ (mm) from 6.7508 to 7.1647 hours UT
DX Back-face corrections:     -14.80 Z4 (nm) or  -0.000561 dZ (mm) from 6.7511 to 7.1650 hours UT

07:20 Stopping the OB and pausing TMS. Injecting spherical (./addsph.py) and take extrafocal pupils: b 72348 and r 72344

07:25 removing z4/z11 and start tmsloop.

07:25 Wait a 3-4 (without nan's) cycles and then slew to high elevation field (elev 86:47), (copointing field for 2013 MW6). Again wait 3-4 cycles here.

07:38 Pause TMS loop, addsph.py and run RB_fastextra4 to take the extra-focal pupil images.

07:40 remsph.py to remove the z4/z11 corrections.

07:40 Taking copointing data and starting tms loop. Copointing.

COPOINTING: B=74209 R=74205
Pointing updates: delta_IE = 14.73", delta_CA = -6.36"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:   -0.08    -0.02    -0.53     1.80
              DX:    0.03    -0.01    -0.22    -0.63

07:45 Observing the field (elev 85 deg) g+r 30-sec integ. The backgrounds are 5500, 3000 ADU and I'll increase it to 60-sec.
  • FWHM ~ 4.7 pix on red and 5.3 pix on blue. DIMM seeing is >1.5"

08:00 I played the script again but this time with the coordinates for 08UT. (Note: the Target Name may be a mess).

r 080923 is trailed (upper left to lower right or vv - not rotationally, diagonally). Next one, r 081232, looks fine.

08:21 FWHM is now better: 3.5/3.2 pix on blue/red. Around this time, I'm noting thermal Z4 ~ 500, 1000 nm and the ambient temp has gone down about 1 deg since we started, to 6 C.
  • My impression is that the blue images do not look as round as the red ones.

robs froze up here and I could not recover it. top

08:46 Stopped OB

08:46 killed the tms_loop.pl (because robs froze) and Heejoo disabled the laser.

08:46 dofpia: First pupils on red are converged, blue pupils are suggesting that a Z4 correction of -898 nm is needed.

08:56 Reconfiguring LUCIS.

Arp 187 - Bino-LUCIs
  • Start with telluric HD 30729
    • putting masks in turnout to check fieldstop alignment
      • L1 0001-0002 - 2.51sec dark
      • L2 0001-0002 - 2.51sec dark
      • L1 0003-0006 - checks
      • L2 0003-0005 - checks
    • UT 09:11 - fieldstops checked
    • UT 09:20 - pointing check
    • UT 09:25 - collimation check
    • UT 09:28 - sending preset
      • seeing in guiders ~0.6"
      • Alignment
      • L1 -10.26", 0.629"
      • L2 -9.60", 0.05"
    • UT 09:35 - starting spectra
    • L1 0010-0012
    • L2 0009-0011
  • Arp 187
    • UT 09:41 - sending preset
    • Guiders showing 0.6"
    • L1 -0.569", 0.510"
    • L2 -0.029", -0.163"
    • L1 +0.12", L2-0.12" tweak
    • UT 09:56 - starting spectroscopy
    • UT 10:14 - seeing 0.7"-0.8"
      • L1 0018-0019, L1 0020-0041, L1 0042-0048
      • L2 0017-0018, L2 0019-0040, L2 0041-0047
    • UT 10:37 - restarting script for another round
    • UT 10:52 - edited script to change -50" dither to -30" - slighlty worried whatever is there at the top of the frame when moving to the bottom of the slit could affect S/N
    • UT 10:56 - seeing worsened a little to 1"-1.2"
    • UT 11:06 - seeing back down to <1"
    • UT 11:17 - staring script for another round (won't finish before hit elevation limit)
    • UT 11:34 - seeing 0.7"-0.8" at airmass 1.919
    • UT 11:39 - completed
  • getting too late to reconfigure to MODS for higher priority science program - will finish off a merger in backup program
  • UGC 4079 Telluric
    • UT 11:40 - sending preset
    • L1 -10.01", -0.090"
    • L2 -9.101", -0.47"
    • L1 +0.12" tweak, L2 ok
    • UT 11:51 - starting spectra
      • L1 0053-0055
      • L2 0051-0053
    • L1 camera wheel failture
      • init camera OK - but double checking
      • -52 N1.8 to 699 for N375, N1.8 again -56, camera wheel unknown again. position is -72 in eng gui at neg limit - move +10 x2, HIRAMO shows 11 on 12 off, initialize, N1.8 is on
    • took L1 exposure manually, continuing with script
  • UGC 4079
    • UT 12:05 - sending preset
    • L1 -0.6397", 0.4005"
    • L2 0.5189", -1.3685"
    • L2 tweak -0.12"
    • UT 12:18 - startng spectra
    • UT 12:58 - as we approach -12deg twilight seeing is 0.4"-0.6"!
      • L1 0060-0067
      • L2 0059-0066
  • UT 13:05 - done, putting LUCIs in safe config
  • UT 13:06 - closing dome
Calibrations:
  • Arp 187 LUCI Cals
    • L1 0068-0085
    • L2 0067-0084
  • UGC 4079 LUCI Cals
    • L1 0086-0103
    • L2 0085-0102
  • MODS Cals
    • 3Kx3K Biases
    • Dual Grating Arcs
    • Slitless Flats
    • Red Imaging Flats SDSS r
UT 13:27 - power bump

MODS Biases finished OK

LUCI cals stopped when the cal unit lamps tried to turn on L1, L2 I paused after arcs off - ARGOS lockout came on.

UT 13:31 - init cal units and that resolved issues

UT 13:31 - continuing Cals LUCI-1

UT 13:36 - continuing MODS Cals

UT 14:05 - LUCI-2 cals done, putting into safe config

UT 14:10 - LUCI-1 cals done putting into safe config

UT 14:15 - putting MODS-2 to bed

UT 14:16 - putting MODS-1 to bed

DONE!

-- BarryRothberg - 12 Nov 2021
Topic revision: r11 - 15 Nov 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback